OpenStudy (anonymous):

y varies jointly with x and the square of n. When y = 25, x = 3 and n = 4. Find the value of y when x = 5 and n = 10. Round to the nearest whole number

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

so the equation is y=xn^2k so you replace the variables into the equation in order to find k 25=3(4)^2k 25=3(16)k 25=48k k=25/48 lets just keep this for when we plug it in since its a large decimal so the equation now is y=25/48xn^2 y=25/48(5)(10^2) y=25/48(5)(100) y=25/48(500) y=260 when rounded to the nearest whole number